OpenStudy (anonymous):

HELPPPPPPP >.< Which of the following functions rises to the right, Equation in comments.

OpenStudy (anonymous):

which one???

OpenStudy (anonymous):

\[u(x)=(1/6)^{x} \] Or \[v(x)=6^{x}\]

hero (hero):

Which do you believe is correct? @laurenbae

OpenStudy (anonymous):

i don't know that's why i asked lol

OpenStudy (anonymous):

if i had to guess maybe the first one...

hero (hero):

Why would you choose the first one? The first one has a "fraction" as its base.

OpenStudy (anonymous):

like i said...it was just a guess...i don't know how to figure this out.

hero (hero):

There are two ways you can figure it out: 1. Algebra 2. Graphing (by hand or electronically)

hero (hero):

To do it using algebra, keep choosing larger and larger values of x, and substitute those values for x into each formula to see what happens. Normally, you would start with values like x = {0,1,2,3} and compare results.

hero (hero):

For example, if we said that x = 3, substituting that into each formula would produce: \(u(3) = (\frac{1}{6})^3 = \frac{1}{216}\) \(v(3) = 6^3 = 216\)

hero (hero):

Clearly, 216 > 1/216

hero (hero):

If you decided to plot both on a graph you would end up with: